# A Stabilized Unfitted Space-time Finite Element Method for Parabolic Problems on Moving Domains

### v1

@@ 4.1 Some auxiliary results
We require the following local trace inequality for cut elements, as stated in [52, Lemma 1].
@@
$\left\|v_{h}\right\|_{K\cap\partial Q}$ $\leq C_{inv,K}h_{K}^{-1/2}\left\|v_{h}\right\|_{K},\quad\forall v_{h}\in\mathbb{P}_{1}(K),$ (4.2)
where $C_{inv,K}$ is a positive constant independent of both the mesh size and the position of the boundary $\partial Q$ relative to the mesh.
@@
where $C_{g,0}$ and $C_{g}$ are positive constants that are independent of the mesh size and the position of the boundary $\partial Q$ relative to the mesh.
@@ 4.2 Stability analysis of the method
$A_{h}(u_{h},v_{h})$ $\lesssim\left\|\thinspace\left|u_{h}\right|\thinspace\right\|_{h,*}\left\|\thinspace\left|v_{h}\right|\thinspace\right\|_{h},\quad\forall u_{h}\in V_{h},~v_{h}\in V_{h},$ (4.6)
@@
$\left(u_{t},v_{h}\right)_{Q}$ $=-\left(u,\partial_{t}v_{h}\right)_{Q}-\left\langle u,v_{h}\right\rangle_{\Sigma_{0}}+\left\langle u,v\right\rangle_{\Sigma_{T}}+\left\langle un_{t},v_{h}\right\rangle_{\Sigma_{s}}$
$\lesssim\|\delta^{-\frac{1}{2}}h^{-1}u\|_{Q}\|\delta^{\frac{1}{2}}h\partial_{t}v_{h}\|_{Q}+\|u\|_{\Sigma_{0}}\|v_{h}\|_{\Sigma_{0}}+\|u\|_{\Sigma_{T}}\|v_{h}\|_{\Sigma_{T}}$
$\quad+\|\gamma^{\frac{1}{2}}h^{-\frac{1}{2}}u\|_{\Sigma_{s}}\|\gamma^{\frac{1}{2}}h^{-\frac{1}{2}}v_{h}\|_{\Sigma_{s}}.$
The fourth term in (3.2) can be bounded using the Cauchy-Schwarz inequality, trace inequality (4.2) and estimate (4.4)
@@
Using integration by parts in time, we obtain
$\left(\partial_{t}v_{h},v_{h}\right)_{Q}=\frac{1}{2}\int_{Q}\partial_{t}(v_{h}^{2})\,\mathrm{d}x=\frac{1}{2}\int_{\partial Q}v_{h}^{2}n_{t}\,\mathrm{d}s=-\frac{1}{2}\|v_{h}\|_{\Sigma_{0}}^{2}+\frac{1}{2}\|v_{h}\|_{\Sigma_{T}}^{2}+\frac{1}{2}\left\langle v_{h}n_{t},v_{h}\right\rangle_{\Sigma_{s}}.$
@@
$\frac{1}{2}\left\langle v_{h}n_{t},v_{h}\right\rangle_{\Sigma_{s}}\leq\frac{1}{2}\|v_{h}\|_{\Sigma_{s}}^{2}=\frac{1}{2}\gamma^{-1}h\|\gamma^{\frac{1}{2}}h^{-\frac{1}{2}}v_{h}\|_{\Sigma_{s}}^{2}\leq C_{0}\gamma^{-1}\|\gamma^{\frac{1}{2}}h^{-\frac{1}{2}}v_{h}\|_{\Sigma_{s}}^{2},$ (4.9)
where $C_{0}=\text{diam}(\widetilde{Q})/2$. To bound the second term, we apply the Cauchy-Schwarz inequality, the $\varepsilon$-inequality $2ab\leq\varepsilon a^{2}+\varepsilon^{-1}b^{2}$, the trace inequality (4.2), and the estimate (4.4)
@@
where $C_{1}=a_{max}^{2}\max_{K\in\mathcal{T}_{h}^{cut}}\left\{C_{inv,K}^{2}h/h_{K}\right\}$ and $C_{2}=C_{1}\max\left\{a_{min}^{-1},1\right\}$. The third term is bounded using the Cauchy-Schwarz inequality and Young’s inequality
@@
where $C_{3}=\text{diam}(\widetilde{Q})^{2}\|\nabla_{x}a\|_{L^{\infty}(Q)}^{2}/2$ and $C_{4}=C_{3}a_{min}^{-1}$. Combining (4.8)–(4.11), we obtain
$A_{h}(v_{h},v_{h})$ $\geq\left(1-C_{2}C_{g}\varepsilon-C_{4}\delta\right)\left(\|a^{\frac{1}{2}}\nabla_{x}v_{h}\|_{Q}^{2}+\left|v_{h}\right|_{g_{h}}^{2}\right)+\left(1-\varepsilon^{-1}\gamma^{-1}-C_{0}\gamma^{-1}\right)\|\gamma^{\frac{1}{2}}h^{-\frac{1}{2}}v_{h}\|_{\Sigma_{s}}^{2}$
@@
where the last inequality follows by choosing $\varepsilon=\left(4C_{2}C_{g}\right)^{-1}$, $\delta\leq\left(4C_{4}\right)^{-1}$, $\gamma\geq 8C_{2}C_{g}+2C_{0}$. ∎
@@ 4.3 Energy error estimates
$\|v-\pi_{h}v\|_{m,Q^{act}_{h}}\lesssim h^{2-m}\left|v\right|_{2,Q^{act}_{h}},\quad\forall 0\leq m\leq 2,\quad\forall v\in H^{2}(Q^{act}_{h}).$ (4.15)
@@
$\|v-\pi_{h}^{e}v\|_{m,Q}=\|Ev-\pi_{h}Ev\|_{m,Q}\leq\|Ev-\pi_{h}Ev\|_{m,Q^{act}_{h}}\lesssim h^{2-m}\|Ev\|_{2,Q^{act}_{h}}\lesssim h^{2-m}\|v\|_{2,Q}.$ (4.16)
@@
$\|v-\pi_{h}^{e}v\|_{m,Q}\lesssim h^{2-m}\|v\|_{2,Q},\quad\forall 0\leq m\leq 2,\quad\forall v\in H^{2}(Q).$ (4.17)
@@
$\mathrm{I}_{2}$ $=\delta\|h\nabla_{x}a\cdot\nabla_{x}\rho\|_{\mathcal{T}_{h}^{act}\cap Q}^{2}+\delta\|ha\nabla_{x}\cdot\nabla_{x}\rho\|_{\mathcal{T}_{h}^{act}\cap Q}^{2}$ (4.22)
$\lesssim\delta h^{2}\|\nabla_{x}a\|_{L^{\infty}(Q)}^{2}\|\nabla\rho\|_{Q}^{2}+\delta h^{2}a_{max}^{2}\left|\rho\right|_{2,Q}^{2}$
@@
$\quad+a_{max}^{2}\left(\|\nabla\rho\|_{Q^{act}_{h}}^{2}+h^{2}\left|\rho\right|_{2,Q^{act}_{h}}^{2}\right)$
@@ 5 Condition number estimates
This section estimates the condition number of the stiffness matrix and shows its independence of the way the boundary $\partial Q$ intersects the mesh. A key ingredient in our analysis is the following space-time Poincaré–Friedrichs inequality:
@@
Assume first that $\Omega_{t}=(a(t),b(t))$ is an interval with $a(t)<b(t)$ for each fixed $t$. By [43, Theorem 3.1] we have
$\left\|w\right\|_{\Omega_{t}}^{2}\leq 2|\Omega_{t}|^{2}\left\|w^{\prime}\right\|_{\Omega_{t}}^{2}+2|\Omega_{t}|(T_{0}w(a))^{2},\quad\forall w\in H^{1}(\Omega_{t}),$ (5.2)
@@
Now suppose $\Omega_{t}$ satisfies Assumption A.1 but is not necessarily an interval. Since $\Omega_{t}$ is a bounded open set in $\mathbb{R}$, it follows from [58, Theorem 1] that $\Omega_{t}$ is a union of countably many disjoint open intervals. Moreover, by Assumption A.1, we have $\left|\partial\Omega_{t}\right|\leq C_{\partial\Omega}$, which implies that $\Omega_{t}$ consists of only finitely many such intervals. Using the bound $\left|\Omega_{t}\right|\leq C_{\Omega}$ together with (5.2), we obtain
$\left\|w\right\|_{\Omega_{t}}^{2}\leq 2C_{\Omega}^{2}\left\|w^{\prime}\right\|_{\Omega_{t}}^{2}+2C_{\Omega}\left\|w\right\|_{\partial\Omega_{t}}^{2},\quad\forall w\in H^{1}(\Omega_{t}).$ (5.3)
@@
By [44, Theorem 1.2], for $d\geq 2$, and exponent $p\in\left[1,d\right)$, if a function $f:\Omega\rightarrow\mathbb{R}$ satisfies $\left\|\nabla f\right\|_{L^{p}(\Omega)},\left\|f\right\|_{L^{p^{\sharp}}(\partial\Omega)}<\infty$, then
@@
where $p^{*}=\frac{dp}{d-p}$, $p^{\sharp}=\frac{(d-1)p}{d-p}$, and the constants $C_{1}(d,p),C_{2}(d,p)$ dependent only on $d$ and $p$.
@@
The result follows directly from [55, Theorem 2.16] combined with the norm equivalence relations (5.10) and (5.11). ∎
@@ 6 Numerical experiments
Based on the error estimate (4.19), the energy-norm error is expected to satisfy $\|a^{\frac{1}{2}}\nabla_{x}(u-u_{h})\|_{Q}\lesssim h\|u\|_{2,Q}$ for any sufficiently smooth solution $u$. To handle numerical integration over curved domains, we apply the approach from [59].
@@ 6.1 Convergence studies for 1D case
where $a_{0}=0.3,b_{0}=0.7$ and $v(t)=\pi\sin(2\pi t)/20$. The finial time is set to $T=1$. This velocity function causes the interval to oscillate periodically over time, as depicted in Figure 3. The time-dependent diffusion coefficient is given by $a(x,t)=0.5t\cos^{2}(x)+0.1$. The source term $f$, boundary data $g$ and initial data $u_{0}$ are prescribed such that the exact solution takes the form
@@ 6.2 Convergence studies for 2D case
with $x_{0}=0.5,y_{0}=0.5,r_{0}=\pi/12$, and the finial time is set as $T=1$. We set the diffusion coefficient $a(x,t)\equiv 1$, and define $f$, $g$, and $u_{0}$ such that the exact solution is given by
@@
where $r=0.7$ and $r_{0}=3.5$. The finial time is set to $T=1$.
@@ 6.3 Small cut effect studies
Varying the shift $l$ alters the relative position between the space-time domain $Q$ and the background mesh $\mathcal{T}_{h}$ potentially leading to small-cut scenarios. We consider $l=\{10^{-5}j\}_{j=0}^{30000}$ and solve the parabolic moving-domain problem (2.1) for each value of $l$, using the same source term $f$, boundary data $g$, initial condition $u_{0}$, and exact solution as in Example 6.2. The parameters are chosen as $\gamma=50$ and $\delta=0.2$. Two stabilization settings are compared: one with ghost penalty stabilization ($\gamma_{1}=0.1$) and one without any stabilization ($\gamma_{1}=0$). For both cases, we compute the condition number $\kappa(A)$ and the relative errors in the $H^{1,0}$- and $L^{2}$-norms. The results are presented in Figure 10.

### v2

@@ 4.1 Some auxiliary results
We require the following local trace inequality for cut elements, as stated in [58, Lemma 1].
@@
$\left\|v_{h}\right\|_{K\cap\partial Q}$ $\leq C_{inv}h_{K}^{-1/2}\left\|v_{h}\right\|_{K},\quad\forall v_{h}\in\mathbb{P}_{1}(K),$ (4.2)
where $C_{inv}$ is a positive constant independent of both the mesh size and the position of the boundary $\Sigma_{s}$ relative to the mesh.
@@
$\left\|\partial_{t}v_{h}\right\|_{Q^{act}_{h}}^{2}$ $\leq\widetilde{C}_{g}\left(\left\|\partial_{t}v_{h}\right\|_{Q}^{2}+\left|v_{h}\right|_{g_{h}}^{2}\right),$ (4.5)
where $C_{g,0},~C_{g}$ and $\widetilde{C}_{g}$ are positive constants that are independent of the mesh size and the position of the boundary $\Sigma_{s}$ relative to the mesh.
@@ 4.2 Stability analysis of the method
$A_{h}(u_{h},v_{h})$ $\lesssim\left\|\thinspace\left|u_{h}\right|\thinspace\right\|_{h,*}\left\|\thinspace\left|v_{h}\right|\thinspace\right\|_{h},\quad\forall u_{h}\in V_{h},~v_{h}\in V_{h}.$ (4.7)
@@
$\left(\partial_{t}u,v_{h}\right)_{Q}$ $=-\left(u,\partial_{t}v_{h}\right)_{Q}-\left\langle u,v_{h}\right\rangle_{\Sigma_{0}}+\left\langle u,v\right\rangle_{\Sigma_{T}}+\left\langle un_{t},v_{h}\right\rangle_{\Sigma_{s}}$
$\leq\ \|\delta^{-\frac{1}{2}}h^{-1}u\|_{Q}\|\delta^{\frac{1}{2}}h\partial_{t}v_{h}\|_{Q}+\|u\|_{\Sigma_{0}}\|v_{h}\|_{\Sigma_{0}}+\|u\|_{\Sigma_{T}}\|v_{h}\|_{\Sigma_{T}}$
$\quad+\gamma^{-1}h\|\gamma^{\frac{1}{2}}h^{-\frac{1}{2}}u\|_{\Sigma_{s}}\|\gamma^{\frac{1}{2}}h^{-\frac{1}{2}}v_{h}\|_{\Sigma_{s}}.$
The fourth term in (3.3) can be bounded using the Cauchy-Schwarz inequality, the trace inequality (4.2) and estimate (4.4)
@@
Without loss of generality, we assume $h<1$. Using integration by parts in time, we obtain
$\left(\partial_{t}v_{h},v_{h}\right)_{Q}=\frac{1}{2}\int_{Q}\partial_{t}(v_{h}^{2})\,\mathrm{d}x\,\mathrm{d}t=\frac{1}{2}\int_{\partial Q}v_{h}^{2}n_{t}\,\mathrm{d}s=-\frac{1}{2}\|v_{h}\|_{\Sigma_{0}}^{2}+\frac{1}{2}\|v_{h}\|_{\Sigma_{T}}^{2}+\frac{1}{2}\left\langle v_{h}n_{t},v_{h}\right\rangle_{\Sigma_{s}}.$
@@
$\frac{1}{2}\left\langle v_{h}n_{t},v_{h}\right\rangle_{\Sigma_{s}}\leq\frac{1}{2}\|v_{h}\|_{\Sigma_{s}}^{2}=\frac{1}{2}\gamma^{-1}h\|\gamma^{\frac{1}{2}}h^{-\frac{1}{2}}v_{h}\|_{\Sigma_{s}}^{2}\leq\frac{1}{2}\gamma^{-1}\|\gamma^{\frac{1}{2}}h^{-\frac{1}{2}}v_{h}\|_{\Sigma_{s}}^{2}.$ (4.10)
To bound the second term, we apply the Cauchy-Schwarz inequality, the $\varepsilon$-inequality $2ab\leq\varepsilon a^{2}+\varepsilon^{-1}b^{2}$, the trace inequality (4.2), the quasi-uniformity (3.1) and the estimate (4.4)
@@
where $C_{1}=a_{max}^{2}C_{inv}^{2}\nu$ and $C_{2}=C_{1}\max\left\{a_{min}^{-1},1\right\}$. The third term is bounded using the Cauchy-Schwarz inequality and Young’s inequality
@@
where $C_{3}=\|\nabla_{x}a\|_{L^{\infty}(Q)}^{2}/2$ and $C_{4}=C_{3}a_{min}^{-1}$. Combining (4.9)–(4.12), we obtain
$A_{h}(v_{h},v_{h})$ $\geq\left(1-C_{2}C_{g}\varepsilon-C_{4}\delta\right)\left(\|a^{\frac{1}{2}}\nabla_{x}v_{h}\|_{Q}^{2}+\left|v_{h}\right|_{g_{h}}^{2}\right)+\left(1-\varepsilon^{-1}\gamma^{-1}-\frac{1}{2}\gamma^{-1}\right)\|\gamma^{\frac{1}{2}}h^{-\frac{1}{2}}v_{h}\|_{\Sigma_{s}}^{2}$
@@
where the last inequality follows by choosing $\varepsilon=\left(4C_{2}C_{g}\right)^{-1}$, $\delta\leq\left(4C_{4}\right)^{-1}$, $\gamma\geq 8C_{2}C_{g}+1$. ∎
@@ 4.3 Energy error estimates
$\|v-\pi_{h}v\|_{m,\mathcal{T}_{h}^{act}}\lesssim h^{2-m}\left|v\right|_{2,Q^{act}_{h}},\quad\forall 0\leq m\leq 2,\quad\forall v\in H^{2}(Q^{act}_{h}).$ (4.16)
@@
$\|v-\pi_{h}^{e}v\|_{m,\mathcal{T}_{h}^{act}\cap Q}$ $=\|Ev-\pi_{h}Ev\|_{m,\mathcal{T}_{h}^{act}\cap Q}\leq\|Ev-\pi_{h}Ev\|_{m,\mathcal{T}_{h}^{act}}$
$\lesssim h^{2-m}\|Ev\|_{2,Q^{act}_{h}}\lesssim h^{2-m}\|v\|_{2,Q}.$
@@
$\|v-\pi_{h}^{e}v\|_{m,\mathcal{T}_{h}^{act}\cap Q}\lesssim h^{2-m}\|v\|_{2,Q},\quad\forall 0\leq m\leq 2,\quad\forall v\in H^{2}(Q).$ (4.17)
@@
$\mathrm{I}_{2}$ $\ \lesssim\ \delta\|h\nabla_{x}a\cdot\nabla_{x}\rho\|_{\mathcal{T}_{h}^{act}\cap Q}^{2}+\delta\|ha\nabla_{x}\cdot\nabla_{x}\rho\|_{\mathcal{T}_{h}^{act}\cap Q}^{2}$ (4.22)
$\lesssim\delta h^{2}\|\nabla_{x}a\|_{L^{\infty}(Q)}^{2}\|\nabla_{x}\rho\|_{Q}^{2}+\delta h^{2}a_{max}^{2}\left\|\rho\right\|_{2,\mathcal{T}_{h}^{act}\cap Q}^{2}$
@@
$\quad+a_{max}^{2}\left(\|\nabla\rho\|_{Q^{act}_{h}}^{2}+h^{2}\|\rho\|_{2,\mathcal{T}_{h}^{act}}^{2}\right)$
@@
Remark 4.2.
The $H^{2}(Q)$ regularity is assumed for convenience in establishing optimal convergence. However, this regularity can be weakened to $H^{1+s}(Q)$ for any $\frac{1}{2}<s\leq 1$. To achieve this estimation, the last term of the definition of $\left\|\thinspace\left|v\right|\thinspace\right\|_{*}^{2}$ is replaced by
$\sum_{K\in\mathcal{T}_{h}^{act}}\left(\|\delta^{\frac{1}{2}}h^{\frac{1}{2}}a\nabla_{x}v\cdot n_{\partial K}\|_{\partial K\cap Q}^{2}+\|\delta^{\frac{1}{2}}h^{\frac{1}{2}}a\nabla_{x}v\cdot n_{x}\|_{K\cap\partial Q}^{2}\right),$
where $n_{\partial K}$ denotes the unit outward normal vector of $\partial K$. Then Theorem 4.3 can be accordingly derived by applying integration by parts to the last term of $a_{h}$ and using (4.5). Following the definition of the fractional Sobolev space [64] and [65, 66], the results from Lemma 4.1, Theorem 4.6 and (4.17) can be extended to fractional order. Proceeding as in Theorem 4.7, the energy error estimate can be generalized to
$\left\|\thinspace\left|u-u_{h}\right|\thinspace\right\|\lesssim h^{s}\|u\|_{1+s,Q},$
provided that the exact solution satisfies $u\in H^{1+s}(Q),~\frac{1}{2}<s\leq 1$.
@@ 5 Condition number estimates
This section estimates the condition number of the stiffness matrix and shows its independence of the way the boundary $\Sigma_{s}$ intersects the mesh. A key ingredient in our analysis is the following space-time Poincaré–Friedrichs inequality:
@@
Assume first that $\Omega_{t}=(a(t),b(t))$ is an interval with $a(t)<b(t)$ for each fixed $t$. By [49, Theorem 3.1] we have
$\left\|w\right\|_{\Omega_{t}}^{2}\leq 2|\Omega_{t}|^{2}\left\|\partial_{x}w\right\|_{\Omega_{t}}^{2}+2|\Omega_{t}|(T_{0}w(a(t)))^{2},\quad\forall w\in H^{1}(\Omega_{t}),$ (5.2)
@@
Now suppose $\Omega_{t}$ satisfies Assumption A.1 but is not necessarily an interval. Since $\Omega_{t}$ is a bounded open set in $\mathbb{R}$, it follows from [67, Theorem 1] that $\Omega_{t}$ is a union of countably many disjoint open intervals. Moreover, by Assumption A.1, we have $\left|\partial\Omega_{t}\right|\leq C_{\partial\Omega}$, which implies that $\Omega_{t}$ consists of only finitely many such intervals. Using the bound $\left|\Omega_{t}\right|\leq C_{\Omega}$ together with (5.2), we obtain
$\left\|w\right\|_{\Omega_{t}}^{2}\leq 2C_{\Omega}^{2}\left\|\partial_{x}w\right\|_{\Omega_{t}}^{2}+2C_{\Omega}\left\|w\right\|_{\partial\Omega_{t}}^{2},\quad\forall w\in H^{1}(\Omega_{t}).$ (5.3)
@@
By [50, Theorem 1.2], for $d\geq 2$, and exponent $p\in\left[1,d\right)$, if a function $f:\Omega\rightarrow\mathbb{R}$ satisfies $\left\|\nabla f\right\|_{L^{p}(\Omega)},\left\|f\right\|_{L^{p^{\sharp}}(\partial\Omega)}<\infty$, then
@@
where $p^{*}=\frac{dp}{d-p}$, $p^{\sharp}=\frac{(d-1)p}{d-p}$, and the constants $C_{1}(d,p),C_{2}(d,p)$ depend only on $d$ and $p$.
@@
Remark 5.1.
Theorem 5.2 imposes only mild assumptions on the domain $\Omega$, requiring merely that it be a bounded Lipschitz domain with boundary of finite Lebesgue measure. Moreover, the generic constant in (5.6) depends only on the dimension $d$ and is independent of whether $\Omega$ is convex or possesses a rough boundary. This follows from the application of the auxiliary result in [50, Theorem 1.2], where the Sobolev inequality (5.7) is shown to hold even for locally Lipschitz domains that are not necessarily bounded.
@@
The result follows directly from [61, Theorem 2.16] combined with the norm equivalence relations (5.10) and (5.11). ∎
@@ 6 Numerical experiments
Based on the error estimate (4.19), the $H^{1,0}$-norm error is expected to satisfy $\|a^{\frac{1}{2}}\nabla_{x}(u-u_{h})\|_{Q}\lesssim h\|u\|_{2,Q}$ for any sufficiently smooth solution $u$.
To handle numerical integration over curved domains, we adopt the approach from [68]. First, a set of sample points is distributed within each element to identify the cut elements. For each cut element, a recursive subdivision algorithm is applied to generate valid subelements for numerical integration. On the uncut subelements, a standard Gaussian quadrature rule is employed for volume integration. For the valid cut subelements, two isoparametric mappings are constructed to define the surface and volume quadrature rules, respectively. We refer to [68] for more details. In our experiments, we employ the quadratic isoparametric mappings and the fourth-order Gaussian quadrature rule for both surface and volume integration to preserve accuracy.
@@ 6.1 Convergence studies for 1D case
where $a_{0}=0.3,b_{0}=0.7$ and $v(t)=\pi\sin(2\pi t)/20$. The final time is set to $T=1$. This velocity function causes the interval to oscillate periodically over time, as depicted in Figure 3. The time-dependent diffusion coefficient is given by $a(x,t)=0.5t\cos^{2}(x)+0.1$. The source term $f$, boundary data $g$ and initial data $u_{0}$ are prescribed such that the exact solution takes the form
@@ 6.2 Convergence studies for 2D case
with $x_{0}=0.5,y_{0}=0.5,r_{0}=\pi/12$, and the final time is set as $T=1$. We set the diffusion coefficient $a(x,t)\equiv 1$, and define $f$, $g$, and $u_{0}$ such that the exact solution is given by
@@
where $r=0.7$ and $r_{0}=3.5$. The final time is set to $T=1$.
@@ 6.3 Small cut effect studies
Varying the shift $l$ alters the relative position between the space-time domain $Q$ and the background mesh $\mathcal{T}_{h}$, potentially leading to small-cut scenarios. We consider $l=\{10^{-5}j\}_{j=0}^{30000}$ and solve the parabolic moving-domain problem (2.1) for each value of $l$, using the same source term $f$, boundary data $g$, initial condition $u_{0}$, and exact solution as in Example 6.2. The parameters are chosen as $\gamma=50$ and $\delta=0.2$. Two stabilization settings are compared: one with ghost penalty stabilization ($\gamma_{1}=0.1$) and one without any stabilization ($\gamma_{1}=0$). For both cases, we compute the condition number $\kappa(A)$ and the relative errors in the $H^{1,0}$- and $L^{2}$-norms. The results are presented in Figure 10.